OpenStudy (anonymous):

rewrite 5/16 + 6/20 with a common denominator

mathslover (mathslover):

\[\large{\frac{5}{16}+\frac{6}{20}}\] Separate them first

mathslover (mathslover):

I mean : \(\large{\frac{5}{16}}\) and \(\large{\frac{6}{20}}\) .. like this

mathslover (mathslover):

\[\large{\frac{5}{16}}\] \[\large{\frac{5\times 20}{16\times 20}}\] \[\large{\frac{100}{320}}\]

mathslover (mathslover):

and similarly for \(\frac{6}{20}\) : \[\large{\frac{6\times 16}{20\times 16}}\] \[\large{\frac{96}{320}}\]

mathslover (mathslover):

got it @b.l.m ?

mathslover (mathslover):

hence we can re unite them : \[\large{\frac{100}{320}+\frac{96}{320}}\]

OpenStudy (anonymous):

im so lost

mathslover (mathslover):

where r u having problem .. ?

mathslover (mathslover):

Ok i will say you an easier way for this

mathslover (mathslover):